OpenStudy (anonymous):

given the function rule f(x)=x^2-3x+2, what is the output of f(-2)

OpenStudy (gorv):

=(-2)^2-3*(-2)+2 =4+6+2 =12

OpenStudy (anonymous):

Thank You for the help Gorv